OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, –2) with a slope of 3/5 ? Answer 5x – 3y = 19 5x – 3y = –19 3x – 5y = –19 3x – 5y = 19

OpenStudy (anonymous):

rewrite in the form y = mx + c - m = slope and then find which fits the (3,-2)

OpenStudy (aroub):

y=mx+b , m=3/5 , (3,-2) -2=3/5(3) +b -2=9/5+b -10/5-9/5 =b -19/2=b so the equation is y=3/5x-19/2

OpenStudy (anonymous):

wich changes to wich one?

OpenStudy (aroub):

oh okay , so 3x – 5y = –19

OpenStudy (aroub):

oh and i meant to write the y-intercept -19/5